OpenStudy (anonymous):

Use the function rule f(x)= 3x-2. Find the output f(1.5). 2.5 –2.5 6.5 –6.5

OpenStudy (anonymous):

\[ f(x)= 3x-2\] \[ f(1.5)= 3(1.5)-2\] can you do it from here?

OpenStudy (anonymous):

all you had to do is plug in 1.5 for x in the function. Now simplify :)

OpenStudy (anonymous):

Yes, great! Thank you so much!(:

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

Um, 2.5?

OpenStudy (anonymous):

\[ f(x)= 3x-2\] \[ f(1.5)= 3(1.5)-2\] \[ f(1.5)= 4.5-2\] \[ f(1.5)= 2.5\]

OpenStudy (anonymous):

yes you are correct :D

OpenStudy (anonymous):

Haha! I /was/ right!

OpenStudy (anonymous):

Thanks again!

OpenStudy (anonymous):

yw :D